# Simulations of gravitational collapse in null coordinates IV: evolving through the event horizon, with an application to the spherical charged scalar field

### v1

@@ IV.4 Singularity excision
As no influence propagates from smaller to larger $x$ in our code for all $x$ with $B<0$, and we are in spherical symmetry, we can use a crude type of similarity excision where we excise gridpoints $i\geq i_{\text{exc}}$, where $i_{\text{exc}}$ is the first grid point where either $R_{i}<0$ or $R_{i}<R_{i-1}/2$. (The second criterion makes sure that the first non-excised point is not too close to $R=0$.) It is essential that we re-determine $i_{\text{exc}}$ after each time we have solved (40) for $R$, that is, in all sub-timesteps. Obviously $i_{\text{exc}}$ is not allowed to decrease from one sub-timestep to the next, or one timestep to the next.
@@ V.1 Charged initial data
so that for $\omega=0$ the scalar field is purely real, and hence for small $\omega$ we expect the charge to be proportional to $\omega$. With a charged scalar field, $q$ sets an inverse length scale in the field equations, so $qx_{c}$, $qd$ and $q/\omega$ are all relevant dimensionless parameters. This also means that changing $q\to\lambda q$ is equivalent to $x_{c}\to x_{c}/\lambda$, $d\to d/\lambda$, and similarly for all other length scales in the initial data and $\omega\to\lambda\omega$.
@@ V.2 Convergence tests
For convergence testing, we have focused on initial data (82) with $x_{c}=0.5$, $d=0.5$, $\omega=2.0$, and the coupling constant $q=1/\sqrt{4\pi}$. We complete the initial data either with $G=1$, and then set $p{\cal A}=0.15$, or with $R=x$, and then set $p{\cal A}=0.2$. These amplitudes are chosen so that $\max_{x}C\simeq 0.5$ in the initial data at $x\simeq 1.2$.
@@
When expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\chi$ to first order, we found that the errors $f_{2}$ and $f_{3}$ of $\hat{\Xi}\psi$ and $\hat{\Xi}\chi$ diverge as $x^{-1}$. This is not in itself a problem, as $h^{2}f_{2}(x_{i})\simeq h^{2}/(ih)=h/i$, so even at $i=1$ the error goes down as $h$. However, we find that by expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\psi$ to second order, we can essentially remove the $h^{2}f_{2}$ error, so that the error is dominated by $h^{3}f_{3}$. This reduces the actual error at $i=1$ by two or three orders of magnitude already at $h=0.04$, and by more at higher resolution.
@@
We note in passing that at $u=0$, where $\psi$, $\chi$ and either $R$ or $G$ or both are still exact, $\Xi\psi$ and $\Xi\psi$ converge to third order.
@@
$E_{xx(2)}$ $:=$ $-4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2})$ (84)
using a different methods from elsewhere in the code. This means in the eR or eG formulations, where $G$ or $R$, respectively, are found by solving $E_{xx}=0$, we have a non-trivial test of the discretization of that process. In the fe formulation, we have a test of the discretization of the solution of $E_{xx}=0$ for either $R$ or $G$ in the initial data, combined with a test of how the subsequent free evolution error leads to increasing violation of the constraint $E_{xx}=0$. We find that $E_{xx(1)}-E_{xx(2)}=0$ converges to second order, except at the boundaries, where it is only first-order. This is because there we use second-order accurate but one-sided finite differences in evaluating the constraint, and therefore in self-convergence testing at the boundaries we compare a one-sided derivatives at the lower resolution with a centred derivative at the higher resolution.
@@ V.4 Type-II critical collapse of a spherical charged scalar field
As our emphasis in this paper is on developing methods for simulating type-II critical collapse on null cones beyond spherical symmetry, rather than on examining charged collapse in spherical symmetry, we examined only two one-parameter families of initial data, namely (82) with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, for either $R(0,x)=x$ or $G(0,x)=1$. We set the coupling constant to $q=1/\sqrt{4\pi}$. During the evolution, we use sdn gauge, given in (60), with the parameter $x_{0}$ adjusted so that $x=x_{0}$ is approximately the past light cone of the accumulation point of echos in near-critical solutions.
@@
As a proxy for the black hole (Hawking) mass and charge, which are strictly speaking defined at future null infinity, we evaluate $M$ and $Q$ on the first symmetry 2-sphere where the Hawking compactness takes the value $C(u,x)=C_{\text{thr}}$. (“First” here means the smallest such $u$.) In the results presented here in the eG formulation, we choose the obvious value $C_{\text{thr}}=1$, but this is not possible in the eR formulation. We discuss the role of the choice of $C_{\text{thr}}$ in Sec. V.6.
Figure 1: Maximum curvature fine-structure for the $G=1$ family of initial data, in the eG formulation: we plot $1/2\ln|\max T|-\gamma\ln(1-p/p_{*})+A$ against $\ln(1-p/p_{*})$. We use the approximate theoretical value $\gamma=0.374$, and the family-dependent offset A=0.23A=-0.23 has been fitted by eye so that the average of the fine-structure is approximately zero at high fine-tuning. The range of $\ln(1-p/p_{*})$ on the horizontal axes has been chosen as in Fig. 1 of Paper II for comparison. (As in Paper II, we have suppressed axis labels to make figures larger.) The thin dotted line has been fitted by eye to the first two peaks from the right and has slope $\delta\gamma=0.012$, meaning that the curvature critical exponent is slightly larger at low fine-tuning.
Figure 2: Black hole mass fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln M-\gamma\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$. $\gamma$ and $A$ and the horizontal range are as for the curvature scaling plot in Fig. 1. Compare also Fig. 2 of Paper II. The thin dotted line has been fitted by eye to the first two peaks from the right. It has slope $\delta\gamma=0.010$, meaning that the mass critical exponent is slightly larger at low fine-tuning, in line with that for that of the curvature to within a plausible fitting error.
@@
The fine-structure of critical scaling on the dispersal side of the threshold of collapse is shown in Fig. 1 for the curvature invariant $T$. On the collapse side, Figs. 2 and 3 show the fine-structure for $M$ and $Q$ of the first (approximate) MOTS. What we actually plot is $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ in Fig. 1, $\ln M-\gamma\ln(1-p/p_{*})+A$ in Fig. 2, and $\ln Q-\mu\ln(1-p/p_{*})+A$ in Fig. 3. The family-dependent offset $A$ has been fitted so the average of $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ is roughly zero at high fine-tuning. $\gamma$ and $\mu$ have been set to their theoretically predicted approximate values $\gamma=0.374$ and $\mu=0.883$ GundlachMartin96 .
@@ V.5 The effect of charge on the critical scaling
At high fine-tuning, the fine-structures are periodic, thus confirming the values of $\gamma$ and $\mu$ from perturbation theory. The fine structure for $T$ has the same amplitude of about $0.15$ and flat top as in Fig. 1 of Paper II, which showed $T$ for a one-parameter family of real, spherically symmetric initial data. At low fine-tuning, the curve has the same periodic structure but rises slightly with $\ln(1-p/p_{*})$. The same observation applies to the periodic fine structure of $M$ at high fine-tuning. It has an amplitude $\simeq 0.03$. From Fig. 3, the fine structure in $\ln Q$ evaluated also appears to be continuous, with amplitude $\simeq 0.15$.
@@
The value of $\mu$ predicted in GundlachMartin96 was verified in collapse simulations in HodPiran97 and Petryk05 , and is verified here again by us. Figures in these last two papers showed a power law with a wiggle, but did not show the wiggle clearly. Hence our Fig. 3 is the first plot of the fine structure of the charge scaling law for any physical system, here the spherical charged scalar field.
@@
At low fine-tuning, we see deviations from the expected scaling laws. Heuristically, these can be described as modifications of the critical exponents $\gamma$ and $\mu$ and the echoing period $\Delta$ (as we did for the critical exponents for the mass and angular momentum in GunBau18 ). To quantify these modifications, we have fitted the thin dotted lines in Figs. 1, 2 and Fig. 3 by eye. These have slopes of $\delta\gamma=0.012$, $\delta\gamma=0.010$, and $\delta\mu=-0.059$, respectively. Note the values of $\delta\gamma$ on the dispersion and collapse sides are roughly the same, as one would expect, but the fitting by eye is somewhat arbitrary. As in GunBau18 , the small change in critical exponent is a convenient quantitative way of describing the fine structure at low fine-tuning, but we do not have a quantitative explanation for it.
@@
For comparison, we have also fine-tuned a family of uncharged $G=1$ initial data, differing from our charged data only by the choice $\omega=0$. We also set ${\cal A}=0.075$ in order to make $p_{*}\simeq 1$. For the eG formulation, we chose $x_{0}=3.525$ and found $p_{*}\simeq 0.95$. For the eR formulation, we chose $x_{0}=3.530$ and found $p_{*}\simeq 0.94$. The fine-structure for $T$ and $M$ (in the eG formulation, with $C_{\text{thr}}=1$) is given in Figs. 4 and 5. Here we can read off more accurately that the amplitude of the fine structure for $M$ is $\simeq 0.032$. As already noted in Paper II, this differs significantly from the values for the amplitude given in the literature. It is remarkable how periodic both fine-structures are already at low fine-tuning, in contrast to the charged case, where $|Q|/M$ has to be quite small before the critical exponents agree with their values calculated in perturbation theory in GundlachMartin96 . That calculation considered charge as a linear perturbation on an uncharged background critical solution (the Choptuik solution). From Figs. 1, 2 and 3, this seems to be accurate for $\ln|1-p/p_{*}|\lesssim-20$. This corresponds to $|Q|/M\lesssim 10^{-4}$.
@@
Here we present results only for a single one-parameter family of charged initial data, which tells us nothing about universality. Both the mass and charge critical exponents appear to agree with their predicted values, however, which is indirect evidence of their universality. We also find that the fine-structure of the mass scaling law looks very similar in the charged and uncharged families. By contrast, we have found initial evidence that the fine-structure of the charge scaling law may not be universal. We will return to this elsewhere.
@@ V.6 The role of the collapse criterion
In the eG and fe formulations we can choose $C_{\text{thr}}=1$, but not in the eR formulation. In Paper II, for example, where we used the eR formulation, we used $C_{\text{thr}}=0.99$ in spherical symmetry and for mildly non-spherical initial data (and $C_{\text{thr}}=0.8$ for the most-nonspherical initial data), and found that the mass fine-structure was discontinuous, including in spherical symmetry.
@@
The exception is the location $x_{C}$ of the first occurrence of $C=1$: it is a very noisy function of $\ln(p/p_{*}-1)$, but with no obvious jumps. The noisiness is explained, in hindsight, by the difficulty of locating the first occurrence of $C=1$: in practice, we find the $x$-location where $C$ goes through 0 at the first timestep where its maximum is larger than 1. Note that $M$ and $Q$ read off at this noisy location are not visibly noisy in Figs. 2 and 3: we believe this is explained by the fact that $M$ and $Q$, as functions of $x$ along the curve $C=1$ quickly asymptote to $M_{\text{BH}}$ and $Q_{\text{BH}}$, and so the values of $M$ and $Q$ read off at $x=x_{C}$ depend only weakly on $x_{C}$.
We can also understand theoretically why the mass fine-structure in spherical symmetry is continuous for $C_{\text{thr}}=1$ but discontinous for $C_{\text{thr}}<1$. We expect the function $C(u,x;p)$ to be continuous in all its arguments. Nevertheless, the set of points $C(u,x;p)=C_{\text{thr}}$ changes discontinously in $p$ whenever a local maximum of $C(u,x)$ rises above $C_{\text{thr}}$, and a new contour line forms. Therefore the location of the first 2-sphere where $C(u,x;p)=C_{\text{thr}}$ also depends discontinuously on $p$. $C_{\text{thr}}=1$ is an exception because the only contour line of $C(u,x;p)=1$ is the apparent horizon, dividing spacetime into a trapped and an untrapped region. But this exception does not apply beyond spherical symmetry, where a 2-sphere with $C=1$ is not in general marginally trapped.
Based on our numerical experiments with different values of $C_{\text{thr}}$, and this theoretical understanding, we now believe that the discontinuity in the mass fine-structure observed in Paper II was an artifact of the choice $C_{\text{thr}}=0.99$, at least in spherical symmetry. Beyond spherical symmetry, we are not sure. We will return to this question elsewhere.
As a practical discovery, we find that plotting the equivalent of Fig. 7 just for the data points obtained in bisection allows us to quickly find the appropriate value of $x_{0}$: if these curves drop to $x=0$ with increased fine-tuning, our choice of $x_{0}$ is too large (and we lose resolution in the self-similar region). If they rise to $x=x_{\text{max}}$ then $x_{0}$ is too small (and the first occurence of the specified threshold value of $C$ actually occurs at larger $x$, outside our numerical domain).

### v2

@@ IV.4 Singularity excision
In the integration of the hierarchy equations from $x=0$, information propagates only from smaller to larger $x$. The same is true for the time evolution equations (74), whereever $B<0$. Our discretisation of the hierarchy and evolution equations respects this causality. In spherical symmetry, we can therefore use a crude type of singularity excision where we excise gridpoints $i\geq i_{\text{exc}}$, where $i_{\text{exc}}$ is the first grid point where either $R_{i}<0$ or $R_{i}<R_{i-1}/2$. (The second criterion makes sure that the first non-excised point is not too close to $R=0$.) It is essential that we re-determine $i_{\text{exc}}$ after each time we have solved (40) for $R$, that is, in all sub-timesteps. Obviously $i_{\text{exc}}$ is not allowed to decrease from one sub-timestep to the next, or from one full timestep to the next.
@@ V.1 Charged initial data
so that for $\omega=0$ the scalar field is purely real, and hence for small $\omega$ we expect the charge to be proportional to $\omega$. With a charged scalar field, $q$ sets an inverse length scale in the field equations, so $qx_{c}$, $qd$ and $q/\omega$ are all relevant dimensionless parameters. This also means that changing $q\to\lambda q$ is equivalent to $x_{c}\to x_{c}/\lambda$, $d\to d/\lambda$, and $\omega\to\lambda\omega$, and similarly for any other length scales in the initial data. Therefore, it may be convenient to fix a length scale in the initial data, such as one of $x_{c}$, $d$ and $\omega$ in (82), and consider $q$ as a parameter of the initial data in its place (rather than a constant of nature). Thus, in one test below, we keep $x_{c}$, $d$ and $\omega$ the same and vary $q$. Varying $q$ has the conceptual advantage that for given initial data $\psi(x)$, $\chi(x)$ and $R(x)$ or $G(x)$, $Q(x)$ in the initial data is manifestly proportional to $q$.
@@ V.2 Convergence tests
For convergence testing, we have focused on the initial data (82) with $x_{c}=0.5$, $d=0.5$, $\omega=2.0$, and the coupling constant $q=1/\sqrt{4\pi}$. We complete the initial data either with $G=1$, and then set $p{\cal A}=0.15$, or with $R=x$, and then set $p{\cal A}=0.2$. These amplitudes are chosen so that $\max_{x}C\simeq 0.5$ in the initial data at $x\simeq 1.2$.
@@
When expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\chi$ to first order, we found that the errors $f_{2}$ and $f_{3}$ of $\hat{\Xi}\psi$ and $\hat{\Xi}\chi$ diverge as $x^{-1}$. This is not in itself a problem, as $h^{2}f_{2}(x_{i})\simeq h^{2}/(ih)=h/i$, so even at $i=1$ the error goes down as $h$. However, we find that by expanding $R\hat{\Xi}\psi$ and $R\hat{\Xi}\chi$ to second order, we can essentially remove the $h^{2}f_{2}$ error, so that the error is dominated by $h^{3}f_{3}$. This reduces the actual error at $i=1$ by two or three orders of magnitude already at $h=0.04$, and by more at higher resolution.
@@
We note in passing that at $u=0$, where $\psi$, $\chi$ and either $R$ or $G$ or both are still exact, $\Xi\psi$ and $\Xi\chi$ converge to third order.
@@
$E_{xx(2)}$ $:=$ $-4\pi R(\psi_{,x}^{2}+\chi_{,x}^{2}),$ (84)
using a different method from elsewhere in the code. This means in the eR or eG formulations, where $G$ or $R$, respectively, are found by solving $E_{xx}=0$, we have a non-trivial test of the discretization of that process. In the fe formulation, we have a test of the discretization of the solution of $E_{xx}=0$ for either $R$ or $G$ in the initial data, combined with a test of how the subsequent free evolution error leads to increasing violation of the constraint $E_{xx}=0$. We find that $E_{xx(1)}-E_{xx(2)}=0$ converges to second order, except at the boundaries, where it is only first-order. This is because there we use second-order accurate but one-sided finite differences in evaluating the constraint, and therefore in self-convergence testing at the boundaries we compare a one-sided derivatives at the lower resolution with a centred derivative at the higher resolution.
@@ V.4 Type-II critical collapse of a spherical charged scalar field
As our emphasis in this paper is on developing methods for simulating type-II critical collapse on null cones beyond spherical symmetry, rather than on charged collapse in spherical symmetry in detail, we examine only four one-parameter families of initial data. In the first of these, $\phi(0,x)$ is given by (82) with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, and we initialise the radial gauge with $G(0,x)=1$. We set the coupling constant to $q=1/\sqrt{4\pi}\simeq 0.2821$. The second family is as the first except that we initialise the radial gauge with $R(0,x)=x$. As already $R_{,x}(0,x)\simeq 1$ in the first family, these two families do not differ qualitatively. The third family is again like the first, except that we set $\omega=0$, so that the scalar field is real and $Q=0$. The fourth family is again like the first, except that we set $q=10/\sqrt{4\pi}\simeq 2.821$, so that $Q(0,x)$ is 10 times as large.
During the evolution, we use sdn gauge, given in (60), with the parameter $x_{0}$ adjusted so that $x=x_{0}$ is approximately the past light cone of the accumulation point of echos in near-critical solutions.
@@
As a proxy for the black hole Hawking mass and charge, which are strictly speaking defined at future null infinity, we evaluate $M$ and $Q$ on the first symmetry 2-sphere where the Hawking compactness takes the value $C(u,x)=C_{\text{thr}}$. (“First” here means the smallest such $u$.) In the results presented here in the eG formulation, we choose the obvious value $C_{\text{thr}}=1$, but this is not possible in the eR formulation. We discuss the role of the choice of $C_{\text{thr}}$ in Sec. V.6.
Figure 1: Maximum curvature fine-structure for the $G=1$ family of initial data, in the eG formulation: we plot $1/2\ln|\max T|-\gamma\ln(1-p/p_{*})+A$ against $\ln(1-p/p_{*})$. We use the approximate theoretical value $\gamma=0.374$, and the family-dependent offset A=0.23A=-0.23 has been fitted by eye so that the average of the fine-structure is approximately zero at high fine-tuning to the threshold of collapse,(that is, in the limit $|p-p_{*}|\to 0$). The range of $\ln(1-p/p_{*})$ on the horizontal axes has been chosen as in Fig. 1 of Paper II for comparison. (As in Paper II, we have suppressed axis labels to make figures larger.) The thin dotted line has been fitted by eye to the first two peaks from the right and has slope $\delta\gamma=0.012$, meaning that the curvature critical exponent is slightly larger at low fine-tuning
Figure 2: Black hole mass fine-structure for the $G=1$ family of initial data, in the eG formulation, with $C_{\text{thr}}=1$: we plot $\ln M-\gamma\ln(p/p_{*}-1)+A$ against $\ln(p/p_{*}-1)$. $\gamma$ and $A$ and the horizontal range are as for the curvature scaling plot in Fig. 1. Compare also Fig. 2 of Paper II. The thin dotted line has been fitted by eye to the first two peaks from the right. It has slope $\delta\gamma=0.010$, meaning that the mass critical exponent is slightly larger at low fine-tuning, in line with that for the curvature, to within a plausible fitting error.
@@
The fine-structure of critical scaling on the dispersal side of the threshold of collapse is shown in Fig. 1 for the curvature invariant $T$. On the collapse side, Figs. 2 and 3 show the fine-structure for $M$ and $Q$ of the first (approximate) marginally outer-trapped surface (from now on, MOTS and FMOTS for the first MOTS). What we actually plot is $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ in Fig. 1, $\ln M-\gamma\ln(1-p/p_{*})+A$ in Fig. 2, and $\ln Q-\mu\ln(1-p/p_{*})+A$ in Fig. 3. The family-dependent offset $A$ has been fitted so the average of $1/2\ln T-\gamma\ln(1-p/p_{*})+A$ is roughly zero at high fine-tuning. $\gamma$ and $\mu$ have been set to their theoretically predicted approximate values $\gamma=0.374$ and $\mu=0.883$ GundlachMartin96 .
@@ V.5 The effect of charge on the critical scaling
At high fine-tuning, the fine-structures are periodic, thus confirming the values of $\gamma$ and $\mu$ from perturbation theory. The fine structure for $T$ has the same amplitude of about $0.15$ and shape as in Fig. 1 of Paper II, which showed $T$ for a one-parameter family of real, spherically symmetric initial data. At low fine-tuning, the curve has the same periodic structure but rises slightly with $\ln(1-p/p_{*})$. The same observation applies to the periodic fine structure of $M$. It has an amplitude $\simeq 0.03$. From Fig. 3, the fine structure in $\ln Q$ evaluated also appears to be continuous, with amplitude $\simeq 0.15$.
@@
The value of $\mu$ predicted in GundlachMartin96 was verified in collapse simulations in HodPiran97 and Petryk05 , and is verified here again by us to be correct sufficiently close to the threshold of collapse. Figures in these last two papers showed a power law with a wiggle, but did not show the wiggle clearly. Hence our Fig. 3 is the first plot of the fine structure of the charge scaling law for any physical system, here the spherical charged scalar field.
@@
At low fine-tuning, we see deviations from the expected scaling laws. Heuristically, these can be described as modifications of the critical exponents $\gamma$ and $\mu$ (as we did for the critical exponents for the mass and angular momentum in GunBau18 ). To quantify these modifications, we have fitted the thin dotted lines in Figs. 1, 2 and Fig. 3 by eye. These have slopes of $\delta\gamma=0.012$, $\delta\gamma=0.010$, and $\delta\mu=-0.059$, respectively. Note the values of $\delta\gamma$ on the dispersion and collapse sides are roughly the same, as one would expect, but the fitting by eye is somewhat arbitrary. As in GunBau18 , the small change in critical exponent is a convenient quantitative way of describing the fine structure at low fine-tuning, but we do not have a quantitative explanation for it.
@@
For comparison, we have also fine-tuned a family of real, and therefore uncharged, initial data with $G=1$, differing from our charged data only by the choice $\omega=0$. We also set ${\cal A}=0.075$ in order to make $p_{*}\simeq 1$. For the eG formulation, we chose $x_{0}=3.525$ and found $p_{*}\simeq 0.95$. For the eR formulation, we chose $x_{0}=3.530$ and found $p_{*}\simeq 0.94$. The fine-structure for $T$ and $M$ (in the eG formulation, with $C_{\text{thr}}=1$) is given in Figs. 4 and 5. Here we can read off more accurately that the amplitude of the fine structure for $M$ is $\simeq 0.032$. As already noted in Paper II, this differs significantly from the values for the amplitude given in the literature. It is remarkable how periodic both fine-structures are already at low fine-tuning, in contrast to the charged case, where $|Q|/M$ has to be quite small before the critical exponents agree with their values calculated in perturbation theory in GundlachMartin96 . That calculation considered charge as a linear perturbation on an uncharged background critical solution (the Choptuik solution). From Figs. 1, 2 and 3, this seems to be accurate for $\ln|1-p/p_{*}|\lesssim-20$. This corresponds to $|Q|/M\lesssim 10^{-4}$ for the black hole.
@@
So far we have presented results only for a single one-parameter family of charged initial data, which tells us nothing about universality. Both the mass and charge critical exponents appear to agree with their predicted values, however, which is indirect evidence of their universality. We also find that the fine-structure of the mass scaling law looks very similar in the charged and uncharged families. By considering a more highly charged family, we have found initial evidence that the fine-structure of the charge scaling law may not be universal.
We consider again complex Gaussian initial data with $x_{c}=1.5$, $d=0.5$, $\omega=2.0$ and ${\cal A}=0.05$, with initial gauge $G(0,x)=1$. The only difference is that we now set $q=10/\sqrt{4\pi}\simeq 2.821$, 10 times the value we investigated before. (As $q$ has dimension of inverse length, this is equivalent to keeping $q$ fixed and changing all other length scales in the numerical domain and initial data by factors of 10.) We find $p_{*}\simeq 1.157$. Once again we find charge scaling roughly with $\mu=0.883$, but the sign of the charge now oscillates in $\ln(p-p_{*})$. The charge fine-structure is shown in Fig. 6. The mass scaling, by contrast, is unchanged (and therefore we do not show it here). We will return to this non-universality elsewhere.
Figure 6: Charge fine-structure for a different family of complex Gaussian initial data (in the eG formulation, with $C_{\text{thr}}=1$.) The initial scalar field profile is the same, and again $G=1$ in the initial data, but $q$ is larger by a factor of 10, compared with the family represented in Fig. 3. For plotting, we set A=0.23A=-0.23, otherwise as in Fig. 3. The charge is positive where the curve is solid and negative where the curve is dashed. (The kinks in $\ln|Q|$ are where $Q$ goes through zero.) This is qualitatively different from Fig. 3, where the charge is always positive.
@@ V.6 The role of the collapse criterion
In the eG and fe formulations we can choose $C_{\text{thr}}=1$, but not in the eR formulation. In Paper II, for example, where we used the eR formulation, we used $C_{\text{thr}}=0.99$ in spherical symmetry and for mildly non-spherical initial data, and $C_{\text{thr}}=0.8$ for the most-nonspherical initial data, and found that the mass fine-structure was discontinuous, including in spherical symmetry.
@@
The exception is the location $x_{C}$ of the first occurrence of $C=1$: it is a very noisy function of $\ln(p/p_{*}-1)$, but with no obvious jumps. The noisiness is explained by the difficulty of locating the first occurrence of $C=1$: in practice, we find the $x$-location where $C$ goes through 0 at the first timestep where its maximum is larger than 1. Note that $M$ and $Q$ read off at this noisy location are not visibly noisy in Figs. 2 and 3: we believe this is explained by the fact that $M$ and $Q$, as functions of $x$ along the curve $C=1$ (the apparent horizon) quickly asymptote to $M_{\text{BH}}$ and $Q_{\text{BH}}$, and so the values of $M$ and $Q$ read off at $x=x_{C}$ depend only weakly on $x_{C}$.
We can understand theoretically why the mass fine-structure in spherical symmetry is continuous for $C_{\text{thr}}=1$ but discontinous for $C_{\text{thr}}<1$. We expect the function $C(u,x;p)$ to be continuous in all its arguments. Nevertheless, the set of points $(u,x)$ where $C(u,x;p)=C_{\text{thr}}$ is in general discontinuous in $p$, a simple example of catastrophe theory. The same is therefore true for the point in this set with the smallest value of $u$.
However, $C_{\text{thr}}=1$ in spherical symmetry is an exception to this generic behaviour, because $C(u,x;p)=1$ is the apparent horizon, which divides spacetime into a trapped and an untrapped region, so it cannot jump in $p$.
Beyond spherical symmetry, a 2-sphere with $C=1$ is not in general marginally trapped, and so does not lie in the apparent horizon. Therefore we generically expect the mass of the first coordinate 2-sphere with $C=1$ to jump in $p$. We will test this when we apply the eG formulation to axisymmetric critical collapse elsewhere.
As a practical discovery, we find that plotting the equivalent of Fig. 8 just for the data points obtained in a bisection to $p_{*}$ allows us to quickly find the appropriate value of $x_{0}$: if these curves drop to $x=0$ with increased fine-tuning, our choice of $x_{0}$ is too large. If they rise to $x=x_{\text{max}}$ then $x_{0}$ is too small. To find $p_{*}$ and $x_{0}$, we then alternate an automated bisection in $p$ (to machine precision) with an adjustment of $x_{0}$ by hand (to about two to four digits).